# Optimizing compilation of error correction codes for $2\times N$ quantum dot arrays and its NP-hardness

### v1

@@ Abstract
The ability to physically move qubits within a register allows the design of hardware-specific error-correction codes which can achieve fault-tolerance while respecting other constraints. In particular, recent advancements have demonstrated the shuttling of electron and hole spin qubits through a quantum dot array with high fidelity. Exploiting this, we design an error correction architecture, consisting merely of two parallel quantum dot arrays, an experimentally validated architecture compatible with classical wiring and control constraints. We develop a suite of heuristic methods for compiling any stabilizer error-correcting code’s syndrome-extraction circuit to run with a minimal number of shuttling operations. In simulation, these heuristics show that fault tolerance can be achieved on several contemporary quantum error-correcting codes requiring only modestly-optimistic noise parameters. Furthermore, we demonstrate how constant column-weight qLDPC codes can be compiled in a provably minimal number of shuttles that scales constantly with code size using Shor-style syndrome extraction. In addition, we provide a proof of the NP hardness of minimizing the number of shuttle operations for codes not in that class.
@@ I Introduction
Motivated by the potential of far easier scalability than other quantum computing architectures due to industrial fabrication compatibility[1] as well as long coherence times[2], in recent years there have been numerous advancements in experimentally realizing spin qubits in electrically defined quantum dots. In particular, the repeated shuttling of an electron or hole up and down an array of quantum dots with a per-hop error rate less than 0.01% [3, 4] has been demonstrated. On this experimentally-validated foundation of being able to physically move qubits along arrays of quantum dots, we design a quantum computing architecture, consisting merely of two of these quantum dot arrays, as shown in Fig. 1. We refer to this setup as a $2\times n$ architecture. Current fabrication restrictions on the fan-out, wiring, and wirebonding of these semiconductor quantum dot arrays make it attractive to explore the possibility of fault-tolerant computation in this configuration of two parallel arrays. Nonetheless, as recent work suggests, this architecture can be used for full fault-tolerant logical computation with the rotated surface code [5]. In this work, we provide techniques for efficiently compiling any quantum error correction code to such a $2\times n$ architecture. In addition, we describe a class of qLDPC codes that always can be compiled to run with a constant number of shuttles, as their property guarantees the existence of a trivial compilation solution under a certain style of fault tolerant syndrome extraction.
@@
While portions of our approach could potentially be applied to running arbitrary circuits on this architecture, we focus our discussion and assumptions on syndrome extraction circuits used in quantum error correction, which is a natural fit for a two-row architecture as we can simply assign data qubits to the top row, and ancillary qubits to the bottom row. Shuttling these two rows past each other enables us to perform gates between distant qubits, enabling the efficient implementation of even qLDPC codes, a very promising class of codes for achieving large scale quantum error correction [6, 7]. This work aims to provide insight on which error correction codes would be most effective for this hardware. Furthermore, we present tools for evaluating how much shuttling error an arbitrary error-correcting stabilizer code would incur.
This paper is structured as follows: In Section II, we present a way to formalize the general minimization of qubit shuttles. In Section III, we then present a simple way of visualizing the problem formalized in the preceding section. This visualization naturally lends itself to simple heuristic methods that, while not always optimal, yield surprisingly effective results. Following that, in Section IV, we discuss a special case of our problem corresponding to fault-tolerant Shor-style syndrome extraction circuits and how this special case gives us enough structure to further improve our compilation. After these discussions of how to reduce the number of shuttling operations, in Section V, through gate-by-gate simulation of the error correction pipeline, we verify that our techniques of $2\times n$ compilation do indeed reduce error rates of running syndrome extraction circuits on this hardware, and show that in certain error regimes our compilation can make the difference of being above or below threshold. This then culminates in section VI, where we discuss how constant column weight LDPC codes compile with a provably minimal number of shuttles under fault tolerant syndrome extraction. To contrast this optimal performance on codes with a certain structure, section VI also discusses how our compilation performs on random codes, and another way of viewing the compilation problem in order to minimize shuttling operations. Finally, in the concluding remarks, we discuss some of the limitations and potential future work. In the Appendix, we provide additional plots of various codes’ performance, a discussion of the cat state preparation required for fault-tolerant syndrome extraction, as well as a proof of the NP-hardness of compilation.
@@ II Foundations of $2\times n$ Circuit Compilation
We formalize the problem as follows: We are given a circuit as an array of $m$ two-qubit gates, $\mathcal{C}=A_{1},\ A_{2},\ ...,\ A_{m}$, and we say that $\delta(A)$ represents the difference in the indices of qubits involved in gate $A$. For example, if $A$ is a CNOT from a qubit at index $i$ to a qubit at index $j$, then $\delta(A)$ = $|i-j|$. We assume we can lay out our qubits such that each gate always involves one qubit from the top row and one qubit from the bottom row (e.g. in a syndrome extraction circuit, the top row would be data qubits and the bottom row would be ancillary qubits). We enumerate the qubits in the $2\times n$ architecture as shown in Fig. 1. We denote the difference in indices between cross-row pairs of qubits at time-step $t$ by $\delta^{t}$, which will be used to represent the relative position of the two rows throughout this paper.
@@
For qubit re-indexing, there are no restrictions on how to define the bijection, as qubit indices are arbitrary in general, while in the case of gate shuffling, gates cannot arbitrarily be reordered. However, within a syndrome extracting circuit for a CSS code, all gates measuring the $X$ stabilizers may be permuted arbitrarily, and the same holds true for gates measuring $Z$ stabilizers. Furthermore, in both CSS and non-CSS codes, there are some permutations between the gates for measuring the $X$ and $Z$ stabilizers that do not affect the functionality of the circuit. Through simulation on the codes we tested (see Section V), we find that assuming all gates can be reordered arbitrarily almost always results in a perfectly functional syndrome extraction circuit after compilation. However if one wanted to guarantee commutation, at least for CSS codes, one could perform syndrome extraction on one type of stabilizers while shuttling in one direction and then extract the other stabilizers while shuttling back in the other direction as proposed in [5]. In this case, one would compile the syndrome extraction circuits for $X$ and $Z$ checks separately (e.g. by re-indexing only ancillary qubits and keeping the data qubit indices fixed).
@@
We refer to a circuit that measures the parity of a weight $w$ stabilizer by applying $w$ gates to a single ancillary qubit in a non-fault tolerant way as “naive syndrome measurement” (see Fig. 3a). For the purposes of compilation, this corresponds to the general problem, as each ancilla may have multiple gates attached to it. We consider only the two-qubit gates in the circuits to be compiled (i.e., single-qubit gates are folded into the neighboring two-qubit gates and not explicitly tracked).
@@ II.1 Gate Shuffling
Now because gates with the same $\delta$ can run concurrently, by merely permuting the gates of $\mathcal{C}$ so that gates of the same $\delta$ are adjacent to one another (sorting $\mathcal{C}$ by $\delta$ is a straightforward way to achieve this), we can partition $\mathcal{C}$ into a collection of sub-arrays for each corresponding different value of $\delta$. The number of such sub-arrays is $|\mathcal{G}|_{0}$, where $|\mathcal{G}|_{0}$ is defined as the number of unique values within $\mathcal{G}$ – the array that contains the $\delta$ values for the current order of gates. All gates in a given sub-array can be executed concurrently, and a single shuttling operation is required to set the architecture in the correct configuration for the corresponding sub-array to be executed. We can therefore state that $|\mathcal{G}|_{0}$ represents the minimum number of required shuttles obtainable by gate shuffling, as we can always trivially sort the gates by their corresponding $\delta$, assuming all gates in $\mathcal{C}$ commute.
Building on this result of how to perform gate shuffling to achieve $|\mathcal{G}|_{0}$ shuttles, we now discuss qubit re-indexing with the goal of finding the bijection of qubit indices that would best minimize $|\mathcal{G}|_{0}$, further reducing the required number of shuttle operations. To solve this, we first fix the data qubits, and only think about re-indexing the ancillary qubits. While the main motivation for fixing the data qubits is to make the problem more tractable, as mentioned before, fixing the data qubits does indeed cause the case of having different $X$ and $Z$ extraction circuits for CSS codes much more manageable as well.
@@ II.2 Reducing Shuttles via Ancillary Qubit Re-indexing
For a given naive syndrome-extraction circuit $\mathcal{C}$, we would like to find a bijection of qubit indices that minimizes the resulting $|\mathcal{G}|_{0}$ (i.e. the number of necessary shuttles). As an example, we refer to Fig. 4, where we would like to reorder qubits 6, 7, and 8, noting that when the index of an ancillary qubit is changed, the $\delta$ values of all gates connected to it will also change. To account for this, we partition $\mathcal{G}$ into $s$ sets, where $s$ is the number of ancillary qubits, and each set will contain the $\delta$ values for the gates touching the corresponding ancillary qubit. In the example present in the left-hand portion of Fig. 4, $\mathcal{G}=[3,4,5,2,3,5]$ is partitioned into sets $R_{6},R_{7},R_{8}$, equal to respectively $\{3,4,5\},\{2,3\},\{5\}$, where $\{3,4,5\}$ are the $\delta$-values for all gates connected to qubit 6, $\{2,3\}$ are the $\delta$ values for all gates connected to qubit 7, and $\{5\}$ is the $\delta$ of the one gate connected to qubit 8. As we are interested in reordering these ancillary qubits, we need to see how their current position affected the values of $\mathcal{G}$. For this reason, we label the ancillary qubits $+1,+2,...,+s$, and notice that each position augments the values in corresponding set element-wise. For example, qubit 6 corresponds to the set $R_{6}=\{3,4,5\}$, which we can then write as $R_{6}=\{2,3,4\}+1$, where the $+1$ encodes that this is the first ancillary qubit. We refer to the set $P_{i}$ of $R_{i}=P_{i}+k_{i}$ as the primitive set of the $i^{\text{th}}$ ancillary qubit ($k_{i}$ is the integer corresponding to the current position of $P_{i}$). The primitive set encodes the relative positions between all top-row (data) qubits that interact with a given bottom-row (ancillary qubit), while $k_{i}$ encodes the current position of that ancillary qubit in the bottom row. On the right-hand side of Fig. 4, the ancillary qubits have been reordered to minimize $|\mathcal{G}|_{0}$. Notice how the primitive sets remain the same under this reordering (how to derive the primitive sets from a given error correcting code is discussed in more detail in Section VI). Furthermore, we can concretely define an admittedly not very friendly combinatorial optimization problem as:
@@ III Heuristics for General Compilation
In order to create the corresponding chain of blocks for a given primitive set $P$, we start with a number of blocks equal to the largest element in $P$ plus one. Then enumerating those blocks from 00 to the largest element in $P$, blocks that correspond to elements in $P$ are solid blocks, and all other blocks are air blocks. The example provided in Fig. 5 has only contiguous solid blocks, however, in general this will not be the case and is a source of difficulty in finding good heuristics (Fig. 11 has an example of non-contiguous solid blocks). The main benefit of this stacking block interpretation is that now it becomes much more simple to reason about heuristics and the problem as a whole. For instance, two rather decent heuristics for finding good orders are:
@@
Because each of these will only take $O(n\log(n))$ time, it is quite efficient to run both and simply take the better of the two choices, breaking ties with the heuristic not being primarily sorted by. Thereupon, as the map from the old order to the new order discovered for packing blocks on the staircase directly corresponds to the bijection $\pi$ mentioned in the problem formulation, we can apply this mapping to the ancillary qubits in the original circuit to a create a new compiled circuit. We call this method ancillary heuristic re-indexing (AHR). Notice that both sorting metrics would work in the simple problem of Fig. 5. To see these heuristics finding optimal solutions on real codes, such as Shor’s 9 qubit code, and Steane’s 7 qubit code, refer to Appendix .1. A third heuristic which is also checked when we mention AHR (the best of three is chosen as the output) is to sort all of the primitive sets by the first element in their set in descending order. Values that have occurred already are deprioritized according to the number of times they’ve occurred. For example, this would sort $\{6,6,6,5,5,4,3,3,2,1,1,0\}$ as $\{6,5,4,3,2,1,0,6,5,3,1,6\}$. This was originally introduced as first guess to solving the case when all primitive sets are size 1 (this corresponds to Shor-syndrome extraction and is discussed in the following section), however, we find that occasionally this method finds a decent solution when the first two heuristics fail. In particular, this third heuristic is the best of the three for the toric code’s naive syndrome extraction circuit.
@@
IV Shor-style fault tolerant Circuit Compilation
In the previous section we discussed the general problem of assigning primitive sets to integers. However, the circuits that the sets were derived from were not quite representative of fault tolerant quantum error correcting circuits, as they do not provide any protection against the propagation of errors to the ancillary qubits. To remedy this and to assess a simplified version of the problem, in this section we consider a special case where all primitive sets are of size 1 (or equivalently, this is the special case where no ancillary qubit will participate in more than a single two-qubit gate), which corresponds directly to fault tolerant Shor-style syndrome extraction [8]. Whereas before we were trying to optimize circuits that use one ancillary qubit per stabilizer check, we now model the circuit that uses $w$ ancillary qubits for a weight $w$ stabilizer check (see Fig. 3b). There are some state-preparation concerns for the ancilla involved in this technique, as discussed in Appendix .3, but they are of no consequence to what we discuss here.
This leaves us with just the optimization of two qubit gates between the data and ancillary qubits as in the previous sections. This cost of extra qubit overhead is balanced not only by guarantees of fault tolerance, but also by additional structure that will allow us to compile the code to run in some cases far fewer shuttles than the naive case. A family of LDPC codes that compile perfectly (it’s trivial to find an optimal reindexing solution for them) with this type of syndrome extraction is discussed in Section VI.

### v2

@@ Abstract
The ability to physically move qubits within a register allows the design of hardware-specific error-correction codes, which can achieve fault-tolerance while respecting other constraints. In particular, recent advancements have demonstrated the shuttling of electron and hole spin qubits through a quantum dot array with high fidelity. It is therefore timely to explore error correction architectures consisting merely of two parallel quantum dot arrays, an experimentally validated architecture compatible with classical wiring and control constraints. Upon such an architecture, we develop a suite of heuristic methods for compiling any Calderbank-Shor-Steane (CSS) error-correcting code’s syndrome-extraction circuit to run with a reduced number of shuttling operations. We demonstrate how column-regular qLDPC codes can be compiled in a provably minimal number of shuttles that is exactly equal to the column weight of the code when Shor-style syndrome extraction is used. We provide tables stating the number of required shuttles for many contemporary codes of interest. In addition, we provide a proof of the NP hardness of minimizing the number of shuttle operations for general codes, even when using Shor syndrome extraction. We also discuss how one could get around this by placing blanks in the ancilla array to achieve minimal shuttles with Shor syndrome extraction on any CSS code, at the cost of longer ancilla arrays.
@@ I Introduction
Motivated by the potential of far easier scalability than other quantum computing architectures due to industrial fabrication compatibility[1] as well as long coherence times[2], in recent years there have been numerous advancements in experimentally realizing spin qubits in electrically defined quantum dots. In particular, the repeated shuttling of an electron or hole up and down an array of quantum dots with a per-hop error rate less than 0.01% [3, 4] has been demonstrated. On this experimentally validated foundation of being able to physically move qubits along arrays of quantum dots, we explore quantum computing architectures consisting merely of two of these quantum dot arrays, as shown in Fig. 1. These hardware layouts have been referred to as a $2\times n$ architectures [5, 6]. Current fabrication restrictions on the fan-out, wiring, and wirebonding of these semiconductor quantum dot arrays make it necessary to explore the possibility of fault-tolerant computation in this restricted configuration of two parallel arrays. Natural implementations of error correction codes for such architectures were recently proposed in [7], however, without yet the capability of optimizing the syndrome extraction circuits under the significant connectivity constraints of the hardware for codes beyond the surface code. In this work, we provide techniques for efficiently compiling any CSS quantum error correction code to such a $2\times n$ architecture. In addition, we show that “column-regular” qLDPC codes can always be compiled to run with a small constant number of shuttles, as their property guarantees the existence of a trivial compilation solution under a certain style of fault-tolerant syndrome extraction.
@@
While portions of our approach could potentially be applied to running arbitrary circuits on this architecture, we focus our discussion and assumptions on syndrome extraction circuits used in quantum error correction, which is a natural fit for a two-row architecture as we can simply assign data qubits to the top row, and ancillary qubits to the bottom row. Shuttling these rows past each other enables us to perform gates between distant qubits, enabling the efficient implementation of codes that even require long-range connectivity such as qLDPC codes, a very promising class of codes for achieving large-scale quantum error correction [8, 9]. This work aims to provide insight on which error correction codes would be most effective for this hardware. Furthermore, we present tools for evaluating how much shuttling error an arbitrary error-correcting stabilizer code would incur. More general layouts, connecting separate $2\times n$ blocks are now also being suggested [10] and our work will be of significant benefit to such large-scale deployments as well.
This paper is structured as follows: In Section II, we present a way to formalize the general minimization of qubit shuttles. In Section III, we then present a simple way of visualizing the problem formalized in the preceding section. This visualization naturally lends itself to simple heuristic methods that, while not always optimal, yield surprisingly effective results for certain circuits. Following that, in Section IV, we discuss a special case of our problem corresponding to fault-tolerant Shor-style syndrome extraction circuits and how this special case gives us enough structure to substantially improve our compilation. After these discussions on how to reduce the number of shuttling operations, in Section V we provide many relevant examples of both column regular and irregular codes, and provide the number of shuttles our methods were able to compile their syndrome extraction circuits to. This then culminates in Section VI, where we discuss how regular column weight LDPC codes compile with a provably minimal number of shuttles under fault-tolerant syndrome extraction. Further elaborating on ways to construct an error correcting scheme with minimum number of shuttle operations of the ancilla array, Section VI also discusses another way of viewing the compilation problem, trading additional space in the ancilla array for the guarantee of minimum shuttling operations on any code. Finally, in the concluding remarks, we discuss some of the limitations and potential future work, such as the importance of investigating post-selected cat state preparation for use with Shor syndrome extraction.
In the Appendix, we provide a proof of the NP-hardness of compilation via reduction from 3-partition, a brief discussion of how the larger error correction pipeline would look including the cat state preparation required for fault-tolerant syndrome extraction on $2\times n$ architecture, as well as a gate-by-gate simulation of the error correction pipeline, verifying that our techniques of $2\times n$ compilation do indeed result in functional circuits syndrome circuits with better error rates corresponding to fewer shuttles. The Appendix also includes a few additional figures and examples to help illustrate the methods presented here.
@@ II Foundations of $2\times n$ Circuit Compilation
We formalize the problem as follows: We are given a circuit as an array of $m$ two-qubit gates, $\mathcal{C}=A_{1},\ A_{2},\ ...,\ A_{m}$, and we say that $\delta(A)$ represents the difference in the indices of qubits involved in gate $A$. For example, if $A$ is a CNOT from a qubit at index $i$ to a qubit at index $j$, then $\delta(A)$ = |ij||i-j|. We assume we can lay out our qubits such that each gate always involves one qubit from the top row and one qubit from the bottom row (e.g. in a syndrome extraction circuit, the top row would be data qubits and the bottom row would be ancillary qubits). Because the way w define $\delta$ only depends on the offsets of the two rows, our abstraction holds when only one of the two rows is allowed to move, and we will assume a stationary row of data qubits throughout this work. A more detailed sketch of our envisioned hardware is presented in Fig. 12. We enumerate the qubits in the $2\times n$ architecture as shown in Fig. 1, and we denote the difference in indices between cross-row pairs of qubits at time-step $t$ by $\delta^{t}$, which will be used to represent the relative position of the two rows throughout this paper.
@@
For qubit re-indexing, there are no restrictions on how to define the bijection, as qubit indices are arbitrary in general, while in the case of gate shuffling, gates cannot arbitrarily be reordered. However, within a syndrome extracting circuit for a CSS code, all gates measuring the $X$ stabilizers may be permuted arbitrarily, and the same holds true for gates measuring $Z$ stabilizers. Similar to the idea of performing syndrome extraction on one type of stabilizers while shuttling in one direction and then extract the other stabilizers while shuttling back in the other as proposed in [7], we also will assume our code is CSS, and compile the syndrome extraction circuits for $X$ and $Z$ checks separately (e.g. by re-indexing only ancillary qubits and keeping the data qubit indices fixed).
@@
We refer to a circuit that measures the parity of a weight $w$ stabilizer by applying $w$ gates to a single ancillary qubit in a non-fault-tolerant way as “naive syndrome measurement” (see Fig. 3a). For the purposes of compilation, this corresponds to the general problem, as each ancilla may have multiple gates attached to it. We consider only the two-qubit gates in the circuits to be compiled (i.e., single-qubit gates are folded into the neighboring two-qubit gates and not explicitly tracked).
@@ II.1 Gate Shuffling
Now because gates with the same $\delta$ can run concurrently, by merely permuting the gates of $\mathcal{C}$ so that gates of the same $\delta$ are adjacent to one another (sorting $\mathcal{C}$ by $\delta$ is a straightforward way to achieve this), we can partition $\mathcal{C}$ into a collection of sub-arrays for each corresponding different value of $\delta$. We call the number of such sub-arrays $|\mathcal{G}|_{0}$, where $|\mathcal{G}|_{0}$ is defined as the number of unique values within $\mathcal{G}$ – the array that contains the $\delta$ values for the current order of gates. All gates in a given sub-array can be executed concurrently, and a single shuttling operation is required to set the architecture in the correct configuration for the corresponding sub-array to be executed. We can therefore state that $|\mathcal{G}|_{0}$ represents the minimum number of required shuttles obtainable by gate shuffling, as we can always trivially sort the gates by their corresponding $\delta$, assuming all gates in $\mathcal{C}$ commute.
Building on this result of how to perform gate shuffling to achieve $|\mathcal{G}|_{0}$ shuttles, we now discuss qubit re-indexing with the goal of finding the bijection of qubit indices that would best minimize $|\mathcal{G}|_{0}$, further reducing the required number of shuttle operations. To solve this, we fix the data qubits, and only think about re-indexing the ancillary qubits. While the main motivation for fixing the data qubits is to make the problem more tractable, as mentioned before, fixing the data qubits also causes the case of having different $X$ and $Z$ extraction circuits for CSS codes more manageable as well.
@@ II.2 Reducing Shuttles via Ancillary Qubit Re-indexing
For a given naive syndrome-extraction circuit $\mathcal{C}$, we would like to find a bijection of qubit indices that minimizes the resulting $|\mathcal{G}|_{0}$ (i.e. the number of necessary shuttles). As an example, we refer to Fig. 4, where we would like to reorder qubits 6, 7, and 8, noting that when the index of an ancillary qubit is changed, the $\delta$ values of all gates connected to it will also change. To account for this, we partition $\mathcal{G}$ into $s$ sets, where $s$ is the number of ancillary qubits, and each set will contain the $\delta$ values for the gates touching the corresponding ancillary qubit. In the example present in the left-hand portion of Fig. 4, $\mathcal{G}=[3,4,5,2,3,5]$ is partitioned into sets $R_{6},R_{7},R_{8}$, equal to respectively $\{3,4,5\},\{2,3\},\{5\}$, where $\{3,4,5\}$ are the $\delta$-values for all gates connected to qubit 6, $\{2,3\}$ are the $\delta$ values for all gates connected to qubit 7, and $\{5\}$ is the $\delta$ of the one gate connected to qubit 8. As we are interested in reordering these ancillary qubits, we need to see how their current position affected the values of $\mathcal{G}$. For this reason, we label the ancillary qubits +1,+2,,+s+1,+2,...,+s, and notice that each position augments the values in corresponding set element-wise. For example, qubit 6 corresponds to the set $R_{6}=\{3,4,5\}$, which we can then write as $R_{6}=\{2,3,4\}+1$, where the +1+1 encodes that this is the first ancillary qubit. We refer to the set $P_{i}$ of $R_{i}=P_{i}+k_{i}$ as the primitive set of the $i^{\text{th}}$ ancillary qubit ($k_{i}$ is the integer corresponding to the current position of $P_{i}$). The primitive set encodes the relative positions between all top-row (data) qubits that interact with a given bottom-row (ancillary qubit), while $k_{i}$ encodes the current position of that ancillary qubit in the bottom row. On the right-hand side of Fig. 4, the ancillary qubits have been reordered to minimize $|\mathcal{G}|_{0}$. Notice how the primitive sets remain the same under this reordering (how to derive the primitive sets from a given error correcting code is discussed in more detail in Section VI). Furthermore, we can concretely define an admittedly not very friendly combinatorial optimization problem as:
@@ III Heuristics for General Compilation
In order to create the corresponding chain of blocks for a given primitive set $P$, we start with a number of blocks equal to the largest element in $P$ plus one. Then enumerating those blocks from 0 to the largest element in $P$, blocks that correspond to elements in $P$ are solid blocks, and all other blocks are air blocks. The example provided in Fig. 5 has only contiguous solid blocks, however, in general this will not be the case and is a source of difficulty in finding good heuristics (Fig. 9 has an example of non-contiguous solid blocks). The main benefit of this stacking block interpretation is that now it becomes much more simple to reason about heuristics and the problem as a whole. For instance, two rather decent initial heuristics for finding good orders are:
@@
Because each of these will only take $O(n\log(n))$ time, it is quite efficient to run both and simply take the better of the two choices, breaking ties with the heuristic not being primarily sorted by. Thereupon, as the map from the old order to the new order discovered for packing blocks on the staircase directly corresponds to the bijection $\pi$ mentioned in the problem formulation, we can apply this mapping to the ancillary qubits in the original circuit to a create a new compiled circuit. We call this method ancillary heuristic re-indexing (AHR). Notice that both sorting metrics would work in the simple problem of Fig. 5. To see these heuristics finding optimal solutions on real (although pedagogical) codes, such as Shor’s 9 qubit code, and Steane’s 7 qubit code, refer to Appendix .1. These techniques also are sufficient to find optimal ancilla orders on column-regular LDPC codes when using Shor syndrome extraction (see Table. 1). A third heuristic which is also checked when we mention AHR (the best of three is chosen as the output) is to sort all of the primitive sets by the first element in their set in descending order. Values that have occurred already are de-prioritized according to the number of times they’ve occurred. For example, this would sort $\{6,6,6,5,5,4,3,3,2,1,1,0\}$ as $\{6,5,4,3,2,1,0,6,5,3,1,6\}$. This was originally introduced as first guess to solving the case when all primitive sets are size 1 (this corresponds to Shor-syndrome extraction and is discussed in the following section), however, we find that occasionally this method finds a decent solution when the first two heuristics fail. For example, this third heuristic is the best of the three for the toric code’s naive syndrome extraction circuit.
@@
IV Shor-style fault-tolerant Circuit Compilation
In the previous section we discussed the general problem of assigning primitive sets to integers. However, the circuits that the sets were derived from were not quite representative of fault-tolerant quantum error correcting circuits, as they do not provide any protection against the propagation of errors to the ancillary qubits. To remedy this and to assess a simplified version of the problem, in this section we consider a special case where all primitive sets are of size 1 (or equivalently, this is the special case where no ancillary qubit will participate in more than a single two-qubit gate), which corresponds directly to fault-tolerant Shor-style syndrome extraction [11]. Whereas before we were trying to optimize circuits that use one ancillary qubit per stabilizer check, we now model the circuit that uses $w$ ancillary qubits for a weight $w$ stabilizer check (see Fig. 3b). There are some state-preparation concerns for the ancilla involved in this technique, as discussed in Appendix .3, but they are of no consequence to what we discuss here.
This leaves us with just the optimization of two qubit gates between the data and ancillary qubits as in the previous sections. This cost of extra qubit overhead is balanced not only by guarantees of fault tolerance, but also by additional structure that will allow us to compile the code to run in some cases far fewer shuttles than the naive case. We show in Section VI that column-regular codes always compile optimally using this syndrome extraction technique.